A firm does not pay a dividend. It is expected to pay its first dividend of $0.25 per share in 3 years (D3). This dividend will grow at 8 percent indefinitely. Using a 10 percent discount rate, compute the value of this stock. a. $2.50 b. $10.33 c. $13.50 d. $13.75 e. $12.50
Answer #1

Cash flows $0.00 $0.00 $0.25 $13.50 | pv @ 10% 0.909091 0.82645 0.75131 0.75131 Present value $0.00 $0.00 $0.19 $10.14 $10.33
Answer #2



Solution :


D0 = D1 = D2 = 0,

D3 = 0.25 ($)

g = 8% = 0.08 = constant for ever from year 3 onwards.

r = 10% = 0.10

=> 1 + r = 1.10


So,


P2 = D2/(r - g)

=> P2 = 0.25/(0.10 - 0.08) = 12.50 ($) 

=> P0 = P2/(1 + r)^2 = 12.50/1.10^2 = 10.33 ($)


Hence, 


Value of the stock = $10.33 : Option b.  (ANSWER)

answered by: Tulsiram Garg